OpenStudy (anonymous):

simplify 10x² -6x³ what this solve

OpenStudy (anonymous):

-2x^2(3x-5) (10-6x)x^2 2(5-3x)X^2

OpenStudy (anonymous):

factor it x^2(10-6x)=0 x^2 = 0 x= 0 and -6x+10 = 0--->subtract 10 -6x = -10---> divide -6 to both sides x = 10/6 ---> simplify x= 5/3

OpenStudy (anonymous):

please don"t factoring

OpenStudy (anonymous):

you're not suppose to factor??

OpenStudy (anonymous):

simplify each expression,if possible .if not possible, write in simplest form

OpenStudy (anonymous):

the expression is simplified, but you can write it as-2x^2(3x-5)

OpenStudy (anonymous):

thank you dear

OpenStudy (anonymous):

you're welcome